what is the exponential function y = ab^x for a graph that includes (1,6) and (0,2)?
Found 2 solutions by stanbon, scott8148:
Answer by stanbon(57250)

what is the exponential function y = ab^x
for a graph that includes (1,6) and (0,2)?
----
(0,2) implies 2 = ab^0
---
(1,6) implies 6 = ab^1
----------
Then: 2 = a
And: 6 = ab
---
So, b = 3
---
Equation:
y = 2*3^x
------------------
Cheers,
Stan H.

Answer by scott8148(6628)

2 = a b^0 ___ 2 = a

6 = 2 b^1 ___ 3 = b

y = 2(3^x)